once upon a time miss abc invited 50 enemies to dinner. solve this question


1
Expert's answer
2021-05-18T14:08:29-0400

Miss ABC can use 0.5 pounds of poison maximum.

Let the mass of X be "x" pounds, then mass of Y is "y=0.5-x" pounds.

The price should be minimised, and the effect should be 50 people killing.


"75x+200(0.5-x)=50"

"75x+100-200x=50"

"125x=50"

"x=\dfrac{50}{125}"

"x=0.4"

Then


"y=0.5-0.4=0.1"

The total price will be 


"0.4(100)+0.1(400)=80(gold\ pieces)"

0.4 pounds X, 0.1 pounds Y, 80 gold pieces.